# Sensing-Assisted Channel Estimation for OFDM ISAC Systems: Framework, Algorithm, and Analysis

### v1

@@ I Introduction
Sensing-assisted channel estimation represents a prominent example of sensing-assisted communication, offering significant advantages for coherent demodulation, especially for orthogonal-frequency-division-multiplexing (OFDM) systems. In an ISAC system, the communication channel estimator can inherently obtain additional prior knowledge from sensing results, which, by intuition, can enhance estimation performance. Therefore, this topic has garnered increasing attention in recent years[12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Some studies consider a monostatic ISAC system. Since sensing and communication channels in the monostatic ISAC system are only partially matched, extra efforts are required to distinguish between communication-relevant paths and irrelevant paths. For example, feedback from users can be leveraged in the channel recovery based on the sparse basis obtained from angular sensing [12], and false path suppression is required before exploiting the delay-Doppler domain sparsity for channel estimation [13]. In [14, 15, 16], channel estimation and target detection are jointly conducted by exploring the sparsity of ISAC channels. However, if a bistatic ISAC system is considered, the sensing and communication channels share identical propagation paths [22]. In [17], the angle of arrival (AoA) obtained by multiple-signal-classification (MUSIC) is utilized for the Kalman-based channel estimation enhancement, while authors in [18] also employ MUSIC for super-resolution angle estimation to aid the channel estimation, reducing training overhead. With assistance from the phased-array radar, authors in [19] turn the channel estimation problem into an angular domain sparse recovery problem. Channel estimation and AoA estimation are jointly refined through angular subspace pruning in [20], and rough estimates of angle and delay are fed into a convolutional neural network (CNN) to enhance channel estimation and parameter estimation in [21]. Most of the existing studies primarily focus on the exploration of angle estimation in assisting channel estimation, while the potential utilization of sensing results related to delay and Doppler remains significantly underexplored.
@@
Beyond the scope of ISAC, parametric model-based channel estimation remains an effective methodology within traditional channel estimation approaches. It involves directly estimating key channel parameters, such as the number of paths, path gain, path delay, and even Doppler shift for each path. Interestingly, the estimation of these parameters aligns well with the principles of sensing-assisted channel estimation, despite the fact that this specific terminology has not been widely adopted in previous research. The parametric model-based channel estimation can be considered as a kind of sensing-assisted channel estimation for bistatic ISAC systems. In the paradigm of parametric model-based channel estimation, the conventional approach involves two sequential steps: first, the estimation of channel parameters (e.g., delay, Doppler, and angle) through signal processing techniques, followed by the estimation of path gains utilizing either a least squares (LS) estimator [23, 24, 25, 26], a linear minimum mean square error (LMMSE) [27, 28] estimator or deep learning methods [29]. Depending on the channel model employed, both the necessary channel parameters and the utilized signal processing techniques may differ. To estimate the multi-path delay, estimation-of-signal-parameters-using-rotational-invariance-technique (ESPRIT) is used in [27, 26] and compressed sensing (CS) is used in [28], while [25] utilizes the Hannan-Quinn (HQ) criterion, which is of reduced complexity. When it comes to multiple-input multiple-output (MIMO) scenarios, angular information is often desired. For example, authors in [23] use MUSIC to estimate both angles and delay, while a subspace-based algorithm is proposed in [29] for angular estimation, and a joint angle and delay estimation is proposed in [24].
@@ III-A Architecture Design
III-B Low Complexity Sensing Algorithm Tailored for the Estimator
@@ III-C Sensing-Assisted Channel Estimation Procedure
The overall procedure of our proposed channel estimation algorithm is shown in Fig. 3. In each slot, the sensing algorithm will feed the coarse sensing results to the channel estimation block. The update check module will determine whether to update the LMMSE coefficients or not. If yes, $\hat{\mathbf{R}}_{\mathbf{h}\mathbf{h}_{\mathrm{p}}}$ and $\hat{\mathbf{R}}_{\mathbf{h}_{\mathrm{p}}\mathbf{h}_{\mathrm{p}}}$ will be updated with sensing results in the current slot. Otherwise, the LMMSE coefficients from the last slot can still be used. Finally, the CFR can be estimated by multiplying the LMMSE channel coefficient matrix by the LS result from the current slot. The ability to tolerate sensing errors and the lowered update frequency is enabled by the tolerance factors incorporated into our algorithm design, as discussed in the next section.
@@ V-A NMSE Analysis
The latter term $b_{i}$ depends on the similarity between the actual and constructed channel covariance matrix $\mathbf{R}_{\mathbf{h}_{\mathrm{p}}\mathbf{h}_{\mathrm{p}}}$ and $\hat{\mathbf{R}}_{\mathbf{h}_{\mathrm{p}}\mathbf{h}_{\mathrm{p}}}$. Denote $\mathbf{D}=\mathbf{V}^{\mathsf{H}}\mathbf{U}$, $b_{i}$ can be further expressed as
@@ VI-A Numerical Results for Tolerance Factors
First, numerical results for evaluating tolerance factors are presented. The NMSE is calculated based on Eq. (7), where $\hat{\sigma}_{w}^{2}$ is set as $10^{-5}$. Three paths with relative SNRs of $[0,-5,-8]$ dB, $\tau=[100,200,400]$ ns, $f_{\mathrm{d}}=[0,-1.87,3.73]$ kHz are considered. Given this scenario, the impact of sensing error on channel estimation performance with fixed tolerance factors is first examined. Then, we vary the size of tolerance factors and observe the channel estimation performance under fixed estimation errors. NMSE performance under different SNRs is evaluated, as shown in Fig. 4.
@@ VI-B Simulation Results for Channel Estimation Performance
Next, the proposed sensing-assisted LMMSE channel estimation method is compared with other channel estimators. Three types of estimators are adopted as the benchmarks: 1) the simplest LS estimator, where LS is applied at pilots with spline interpolation for data symbols; 2) the robust LMMSE estimator, where the maximum delay spread and Doppler spread are assumed to be perfectly achieved and uniform distributions are adopted for the construction of channel correlation matrix; 3) the parametric model-based estimator, where the high-resolution MUSIC algorithm is applied for the estimation of delay and Doppler, and the LMMSE algorithm is further conducted for the estimation of the amplitude of each path, which is an extended method of [27] for the doubly-selective channel. In the parametric model-based method, the accuracy of delay and Doppler estimation depends on the step size in the MUSIC spectrum search. Two cases are considered: 1) the same size as our 2D FFT bin; and 2) a finer estimation with a step size of only 1/4 of the FFT bin for both delay and Doppler estimation. A small step size leads to high accuracy, but highly increases the computational complexity in MUSIC and also the estimation of the amplitude of each path. The sensing algorithm uses $S=10$ to reduce Doppler resolution. The tolerance factors are assigned with the sensing resolution values. The simulation considers two consequent slots. Moreover, to validate the effectiveness of applying a constant multi-path intensity profile in the construction of our channel correlation matrices, i.e., $\theta(\tau_{l})=1$ for $l=1,2,\cdots,\hat{L}$, the results with true multi-path intensity profile (perfect gain) are also provided.
@@
(a) NMSE performance for different pilot subcarrier intervals.
(b) NMSE performance for different pilot symbol intervals.
@@
Finally, the impact of the pilot interval on the proposed channel estimator is studied. Following the 3-path model used in the previous simulations, the NMSE performance under different DMRS deployments is shown in Fig. 8. A high SNR of 40 dB is considered in this case study, which can support a low BER for high-order modulations. The robust LMMSE estimator is comparable to the proposed estimator, since the robust LMMSE estimator performs better than other benchmarks in high SNR regions. In Fig. 8(a), the symbol interval is set to 8, i.e., $\Delta_{\text{sym}}=8$, and the impact of different subcarrier intervals is evaluated. The proposed estimator outperforms the robust LMMSE estimator for all subcarrier intervals. The performance of our proposed scheme with the subcarrier interval of $\Delta_{\text{sc}}=10$ is still much better than the robust LMMSE scheme with $\Delta_{\text{sc}}=4$. In Fig. 8(b), the pilot subcarrier interval is set fixed, i.e., $\Delta_{\text{sc}}=8$, and the impact of different pilot symbol intervals is evaluated. We can see that our proposed estimator still performs better than the robust LMMSE estimator for all the pilot symbol intervals. Both cases show that with our proposed scheme, the overhead of pilots can be highly reduced, due to the high efficiency of our proposed scheme.

### v2

# Sensing-Assisted Channel Estimation for Bistatic OFDM ISAC Systems: Framework, Algorithm, and Analysis

@@ I Introduction
Sensing-assisted channel estimation represents a prominent example of sensing-assisted communication, offering significant advantages for coherent demodulation, especially for orthogonal-frequency-division-multiplexing (OFDM) systems. In an ISAC system, the communication channel estimator can inherently obtain additional prior knowledge from sensing results, which, by intuition, can enhance estimation performance. Therefore, this topic has garnered increasing attention in recent years[12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Some studies consider a monostatic ISAC system. Since sensing and communication channels in the monostatic ISAC system are only partially matched, extra efforts are required to distinguish between communication-relevant paths and irrelevant paths. For example, feedback from users can be leveraged in the channel recovery based on the sparse basis obtained from angular sensing [12], and false path suppression is required before exploiting the delay-Doppler domain sparsity for channel estimation [13]. In [14, 15, 16], channel estimation and target detection are jointly conducted by exploring the sparsity of ISAC channels. However, if a bistatic ISAC system is considered, the sensing and communication channels share identical propagation paths [23]. In [17], the angle of arrival (AoA) obtained by multiple-signal-classification (MUSIC) is utilized for the Kalman-based channel estimation enhancement, while authors in [18] also employ MUSIC for super-resolution angle estimation to aid the channel estimation, reducing training overhead. With assistance from the phased-array radar, authors in [19] turn the channel estimation problem into an angular domain sparse recovery problem. Channel estimation and AoA estimation are jointly refined through angular subspace pruning in [20], and rough estimates of angle and delay are fed into a convolutional neural network (CNN) to enhance channel estimation and parameter estimation in [21]. In [22], a MIMO radar is used to perform subspace reconstruction for angle estimation, which assists the following deep learning based gain estimation. Most of the existing studies primarily focus on the exploration of angle estimation in assisting channel estimation, while the potential utilization of sensing results related to delay and Doppler remains significantly underexplored.
@@
Beyond the scope of ISAC, parametric model-based channel estimation remains an effective methodology within traditional channel estimation approaches. It involves directly estimating key channel parameters, such as the number of paths, path gain, path delay, and even Doppler shift for each path. Interestingly, the estimation of these parameters aligns well with the principles of sensing-assisted channel estimation, despite the fact that this specific terminology has not been widely adopted in previous research. The parametric model-based channel estimation can be considered as a kind of sensing-assisted channel estimation for bistatic ISAC systems. In the paradigm of parametric model-based channel estimation, the conventional approach involves two sequential steps: first, the estimation of channel parameters (e.g., delay, Doppler, and angle) through signal processing techniques, followed by the estimation of path gains utilizing either a least squares (LS) estimator [24, 25, 26, 27, 28], a linear minimum mean square error (LMMSE) [29, 30] estimator or deep learning methods [22]. Depending on the channel model employed, the channel parameters and utilized signal processing techniques may differ. To estimate the multi-path delay, estimation-of-signal-parameters-using-rotational-invariance-technique (ESPRIT) is used in [29, 27] and compressed sensing (CS) is used in [30], while [26] utilizes the Hannan-Quinn (HQ) criterion, which is of reduced complexity. When it comes to multiple-input multiple-output (MIMO) scenarios, angular information is often desired. Many methods have been developed for estimating channel parameters including angles [24, 25, 28].
@@ III-A Architecture Design
III-B Low-Complexity Sensing Algorithm Tailored for the Estimator
@@ III-C Sensing-Assisted Channel Estimation Procedure
The overall procedure of our proposed channel estimation algorithm is shown in Fig. 3. In each slot, the sensing algorithm will feed the coarse sensing results to the channel estimation block. The update check module will determine whether to update the LMMSE coefficients according to rules stated in Section IV-C. If yes, $\hat{\mathbf{R}}_{\mathbf{h}\mathbf{h}_{\mathrm{p}}}$ and $\hat{\mathbf{R}}_{\mathbf{h}_{\mathrm{p}}\mathbf{h}_{\mathrm{p}}}$ will be updated with sensing results in the current slot. Otherwise, the LMMSE coefficients from the last slot can still be used. Finally, the CFR can be estimated by multiplying the LMMSE channel coefficient matrix by the LS result from the current slot. The ability to tolerate sensing errors, which can also support a low update rate, is enabled by the tolerance factors incorporated into our algorithm design, as discussed in the next section.
@@ V-A NMSE Analysis
The latter term $b_{i}$ depends on the similarity between the actual and constructed channel correlation matrix $\mathbf{R}_{\mathbf{h}_{\mathrm{p}}\mathbf{h}_{\mathrm{p}}}$ and $\hat{\mathbf{R}}_{\mathbf{h}_{\mathrm{p}}\mathbf{h}_{\mathrm{p}}}$. Denote $\mathbf{D}=\mathbf{V}^{\mathsf{H}}\mathbf{U}$, $b_{i}$ can be further expressed as
@@ VI-A Numerical Results for Tolerance Factors
First, numerical results for evaluating tolerance factors are presented. The NMSE is calculated based on Eq. (7), where $\hat{\sigma}_{w}^{2}$ is set as $10^{-5}$. Three paths with relative SNRs of $[0,-5,-8]$ dB, $\tau=[100,200,400]$ ns, $f_{\mathrm{d}}=[0,-1.87,3.73]$ kHz are considered. Given this scenario, the impact of sensing error on channel estimation performance with fixed tolerance factors is first examined. Then, we vary the size of tolerance factors and observe the channel estimation performance under fixed sensing errors. NMSE performance under different SNRs is evaluated, as shown in Fig. 4.
@@ VI-B Simulation Results for Channel Estimation Performance
Next, the proposed sensing-assisted LMMSE channel estimation method is compared with other channel estimators. Three types of estimators are adopted as the benchmarks: 1) the simplest LS estimator, where LS is applied at pilots with spline interpolation for data symbols; 2) the robust LMMSE estimator, where the maximum delay spread and Doppler spread are assumed to be perfectly achieved and uniform distributions are adopted for the construction of channel correlation matrix; 3) the parametric model-based estimator, where the high-resolution MUSIC algorithm is applied for the estimation of delay and Doppler, and the LMMSE algorithm is further conducted for the estimation of the amplitude of each path, which is an extended method of [29] for the doubly-selective channel. In the parametric model-based method, the accuracy of delay and Doppler estimation depends on the step size in the MUSIC spectrum search. Two cases are considered: 1) the same size as our 2D FFT bin; and 2) a finer estimation with a step size of only 1/4 of the FFT bin for both delay and Doppler estimation. A small step size leads to high accuracy, but highly increases the computational complexity of the adopted sensing algorithm. The sensing algorithm uses $S=10$ to reduce Doppler resolution. The tolerance factors are assigned with the sensing resolution values. The simulation considers two consequent slots. Moreover, to validate the effectiveness of applying a constant multi-path intensity profile in the construction of our channel correlation matrices, i.e., $\theta(\tau_{l})=1$ for $l=1,2,\cdots,\hat{L}$, the results with true multi-path intensity profile (perfect gain) are also provided.
@@
(a) Impact of different pilot subcarrier intervals.
(b) Impact of different pilot symbol intervals.
@@
Finally, the impact of the pilot interval on the proposed channel estimator is studied. Following the 3-path model used in the previous simulations, the NMSE performance under different DMRS deployments is shown in Fig. 8. An SNR of 40 dB is considered in this case study, which can support a low BER for high-order modulations. The robust LMMSE estimator is comparable to the proposed estimator, since the robust LMMSE estimator performs better than other benchmarks in high SNR regions. In Fig. 8(a), the symbol interval is set to 8, i.e., $\Delta_{\text{sym}}=8$, and the impact of different subcarrier intervals is evaluated. The proposed estimator outperforms the robust LMMSE estimator for all subcarrier intervals. The performance of our proposed scheme with the subcarrier interval of $\Delta_{\text{sc}}=10$ is still much better than the robust LMMSE scheme with $\Delta_{\text{sc}}=4$. In Fig. 8(b), the pilot subcarrier interval is set fixed, i.e., $\Delta_{\text{sc}}=8$, and the impact of different pilot symbol intervals is evaluated. We can see that our proposed estimator still performs better than the robust LMMSE estimator for all the pilot symbol intervals. Both cases show that with our proposed scheme, the overhead of pilots can be highly reduced, due to the high efficiency of our proposed scheme.